# Asymmetric dyonic multi-centered rotating black holes

### v1

@@ II The Teo-Wan solutions
where the function $\phi$, the component $A_{\mu}$ of the $1$-form ${\bm{A}}=A_{\mu}dx^{\mu}$ and the 4D metric $g_{\mu\nu}$ ($\mu,\nu=0,\ldots,3$) do not depend on the fifth spatial coordinate $x^{5}$ which has a period of $2\pi R_{KK}$. As is well-known, the dimensional reduction of the 5D Einstein theory leads the 4D Einstein-Maxwell-dilaton theory, described by the action
@@ III.2 $SL(3,{\mathbb{R}})$ nonlinear sigma model
$S=\int\left(R^{h}-G_{AB}\frac{\partial\Phi^{A}}{\partial x^{i}}\frac{\partial\Phi^{B}}{\partial x^{j}}h^{ij}\right)\sqrt{|h|}d^{3}x\,$, ($2$)
@@ III.4 Asymptotically flat solutions of Clément
where $f$ and $g$ are harmonic functions on ${\mathbb{E}}^{3}$, $\eta$ and $A$ are $3\times 3$ constant matrices. To ensure that the 4D metric $ds^{2}_{(4)}=g_{\mu\nu}dx^{\mu}dx^{\nu}$ is asymptotically flat, the matrix $\eta$ must be chosen as
@@
if the two harmonic functions vanish at infinity. If the matrix $A$ satisfies
@@ IV.1 The slowly rotating extremal limit of the Rasheed-Larsen solution
This slowly rotating extremal solution belongs to the class of solutions (36) depending on two harmonic functions of Clément, where the corresponding harmonic functions $f$, $g$ and the matrix $A$ are given by
@@ V.1 Near-horizon geometry
The 4D metric apparently diverges at the points ${\bm{x}}={\bm{x}}_{i}$ but we show that they correspond to smooth Killing horizons, provided the slowly rotating conditions $|J_{i}|<|P_{i}Q_{i}|$ for all $i$. Introducing $r:=|{\bm{x}}-{\bm{x}}_{i}|$, and the standard spherical coordinates $(x,y,z)=(r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta)$, we consider the limit of $r\to 0$. In this limit, the 4D metric, Maxwell gauge potential and scalar field at $r\to 0$ behaves as
@@ V.2 Asymptotic structure
In terms of the standard spherical coordinates $(x,y,z)=(r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta)$, the functions, $f,g,H_{\pm}$ and the $1$-forms ${\bm{\omega}}^{0}$, $\tilde{\bm{\omega}}^{5}$ behave asymptotically as at $r\to\infty$

### v3

@@ II The Teo-Wan solutions
where the function $\phi$, the component $A_{\mu}$ of the 11-form ${\bm{A}}=A_{\mu}dx^{\mu}$ and the 4D metric $g_{\mu\nu}$ ($\mu,\nu=0,\ldots,3$) do not depend on the fifth spatial coordinate $x^{5}$ which has a period of $2\pi R_{KK}$. As is well-known, the dimensional reduction of the 5D Einstein theory leads the 4D Einstein-Maxwell-dilaton theory, described by the action
@@ III.2 $SL(3,{\mathbb{R}})$ nonlinear sigma model
$S=\int\left(R^{h}-G_{AB}\frac{\partial\Phi^{A}}{\partial x^{i}}\frac{\partial\Phi^{B}}{\partial x^{j}}h^{ij}\right)\sqrt{|h|}d^{3}x\,$, (22)
@@ III.4 Asymptotically flat solutions of Clément
where $f$ and $g$ are harmonic functions on ${\mathbb{E}}^{3}$, $\eta$ and AA are $3\times 3$ constant matrices. To ensure that the 4D metric $ds^{2}_{(4)}=g_{\mu\nu}dx^{\mu}dx^{\nu}$ is asymptotically flat, the matrix $\eta$ must be chosen as
@@
if the two harmonic functions vanish at infinity. If the matrix AA satisfies
@@ IV.1 The slowly rotating extremal limit of the Rasheed-Larsen solution
This slowly rotating extremal solution belongs to the class of solutions (36) depending on two harmonic functions of Clément, where the corresponding harmonic functions $f$, $g$ and the matrix AA are given by
@@ V.1 Near-horizon geometry
The 4D metric apparently diverges at the points ${\bm{x}}={\bm{x}}_{i}$ but we show that they correspond to smooth Killing horizons, provided the slowly rotating conditions $|J_{i}|<|P_{i}Q_{i}|$ for all ii. Introducing $r:=|{\bm{x}}-{\bm{x}}_{i}|$, and the standard spherical coordinates $(x,y,z)=(r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta)$, we consider the limit of $r\to 0$. In this limit, the 4D metric, Maxwell gauge potential and scalar field at $r\to 0$ behaves as
@@ V.2 Asymptotic structure
In terms of the standard spherical coordinates $(x,y,z)=(r\sin\theta\cos\phi,r\sin\theta\sin\phi,r\cos\theta)$, the functions, $f,g,H_{\pm}$ and the 11-forms ${\bm{\omega}}^{0}$, $\tilde{\bm{\omega}}^{5}$ behave asymptotically as at $r\to\infty$